# USTC-TD: A Test Dataset and Benchmark for Image and Video Coding in 2020s

### v2

@@ VI Conclusion
In this paper, we propose a test dataset (named USTC-TD) for compression-related research, which covers more diverse content factors. To evaluate the efficiency of USTC-TD, we qualitatively evaluate the USTC-TD on different image/video features and compare it with the previous image/video common test datasets to verify its wide coverage. In addition, we evaluate the advanced compression methods under different metrics benchmarked on USTC-TD, and further analyze their performance to point out the inspirations for future compression-related research.
@@
In the present dataset construction process, we only consider the basic image and video test datasets. In the future, we plan to progressively extend the annotation datasets of USTC-TD for image/video coding for machine (ICM/VCM)[65, 66, 67], such as object segmentation [68], object detection [69], action recognition [70], et al, and the reconstruction dataset of video enhancement, such as image/video super-resolution[71, 72], denoising[73], warping[74], et al, for the testing of compression-related downstream researches, and provide a comprehensive baseline to promote the development of compression-related diverse tasks.
VII Copyright
The released images and sequences are captured and processed by the University of Science and Technology of China (USTC). All intellectual property rights remain with USTC.
The following uses are allowed for the contributed dataset: (1) Data (images and videos) may be published in research papers, technical reports, and development events. (2) Data (images and videos) may be utilized for standardization activities (e.g., ITU, MPEG, AVS, VQEG).
The following uses are NOT allowed for the contributed dataset: (1) Do not publish snapshots in product brochures. (2) Do not use video for marketing purposes. (3) Redistribution is not permitted. (4) Do not use it in television shows, commercials, or movies.

### v3

@@ VII Conclusion
In this paper, we propose a test dataset (named USTC-TD) for compression-related research, which covers more diverse content factors. To evaluate the efficiency of USTC-TD, we qualitatively evaluate the USTC-TD on different image/video features and compare it with the previous image/video common test datasets to verify its excellent compensation. In addition, we evaluate the advanced compression schemes under different metrics benchmarked on USTC-TD, and further analyze their performance to point out the inspirations for future compression-related research.
@@
In the present dataset construction process, we only consider the basic image and video test datasets. In the future, we plan to progressively extend the annotation datasets of USTC-TD for image/video coding for machine (ICM/VCM)[71, 72, 73], such as object segmentation [74], object detection [75], action recognition [76], et al, and the reconstruction dataset of video enhancement, such as image/video super-resolution[77, 78], denoising[79], et al, for the testing of compression-related downstream researches, and provide a comprehensive baseline to promote the development of compression-related diverse tasks.